OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

Parth (parthkohli):

The point slope form, maybe?\[y - y_0 = m(x - x_0)\]Where \((x_0, y_0)\) is your point and \(m\) is your slope.

OpenStudy (anonymous):

Hmm? Can you show me how you would do that?

Parth (parthkohli):

Here, you have \(x_0 = 2\), \(y_0 = 5\) and \(m = 2\). Can you substitute those values in \(y - y_0 = m(x - x_0)\)?

OpenStudy (anonymous):

So, y - 5 = 2(x - 2)?

Parth (parthkohli):

That's right. Simplify that equation by bringing it into the form \(y = \cdots\)

OpenStudy (anonymous):

Ok, Im a little confused on the first part would we subsitute 1 for y and do 1-5?

Parth (parthkohli):

What's the first part?

OpenStudy (anonymous):

y -5

Parth (parthkohli):

No, you would leave \(x\) and \(y\) the way they already are. :-)

Parth (parthkohli):

You have \(y - 5 = 2(x - 2)\). First expand \(2(x - 2)\).

OpenStudy (anonymous):

Ok , 2x - 4

Parth (parthkohli):

Yes, now bring that \(-5\) to the other side.

OpenStudy (anonymous):

By adding?

Parth (parthkohli):

I mean... you have \(y -5 = 2x - 4\). And you have to get \(y\) by itself, so...?

Parth (parthkohli):

Yesh, you add \(5\) to both sides!

OpenStudy (anonymous):

So y = 2x - 9?

Parth (parthkohli):

Hmm...

Parth (parthkohli):

\[y = 2x - 4 + 5\]

Parth (parthkohli):

What is \(-4 + 5\)?

OpenStudy (anonymous):

OHHHH! -4

Parth (parthkohli):

Hmm...

OpenStudy (anonymous):

y = 2x + 1

Parth (parthkohli):

Correct!

OpenStudy (anonymous):

Yay! Thank you soo much!

Parth (parthkohli):

No problem :-)